# supplementary

### v1

@@ 4.5 Explainability of generative AI voting
The consistency of the individual AI choices (see Section 4.3) is modeled as the dependent variable in a predictive machine learning framework. Causal relationships that explain how personal human traits (independent variables) contribute to consistency are studied. We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$. We also assess a different prediction model with dependent variable the top single choice, which is the most preferred alternative, see Figure LABEL:fig:explaination-top-choice for more details. Moreover, further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set.
Explainability of choices. There is no established ground truth of how personal human traits (features) influence voter ballots, and thus we introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on individual choice consistency. It consists of the following methods: (i) Category gradient boosting [70] via which the relative importance of each personal human trait (feature) is derived. Results are shown in Figure 5. (ii) Feature ablation studies [71] via which a specific personal human trait (feature) is removed to understand the impact on the classification performance. The significance of the difference in the prediction score with and without the feature is calculated and comaprisons are made across each single removed feature  [72, 70]). Results are shown in Figure LABEL:fig:explaination-consistency1. (iii) Feature permutation methods [73], wherein feature data points are shuffled to measure the impact on the classification performance (Table LABEL:table:error).
@@ Acknowledgements
Preprints
This manuscript was posted on a preprint: https://doi.org/

### v2

@@ 4.5 Explainability of generative AI voting
The consistency of the individual AI choices (see Section 4.3) is modeled as the dependent variable in a predictive machine learning framework. Causal relationships that explain how personal human traits (independent variables) contribute to consistency are studied. We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$. Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set. We also assess prediction models with different dependent variables, see Figure LABEL:fig:explaination-top-choice for more details.
Explainability of choices. We introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on the consistency of individual choice. We use model agnostic Shapley Additive Explanations [70] to extract individual contributions of each trait. Results are shown in Figure 5. Feature ablation study [71] is used to calculate the error (loss) in the overall prediction accuracy of the model when a feature is removed (results in Table LABEL:table:error).
@@ Acknowledgements
Data Availability
The data for US elections are available in the link https://figshare.com/collections/Generative_AI_Voting_-_ANES/7261288. The data regarding the campaign in Aarau is available on request.